OpenStudy (anonymous):

What is the simplified form of x+3/x^2-x-12 times x-4/x^2-8x+16?

jhonyy9 (jhonyy9):

(x+3) (x-4) ---------- * -------- = (x-3)(x+4) (x-4)^2 so hence are these equations factorized than now can you simplifie ,can you continue ?

jhonyy9 (jhonyy9):

sorry the first fraction not is right so here is correct (x+3) (x-4) ---------- * ---------- = (x+3)(x-4) (x-4)^2

jhonyy9 (jhonyy9):

Tiffani can you ending ?

jhonyy9 (jhonyy9):

so on the first fraction simplifie by (x+3) and will remained 1/(x-4) in the second fraction simplifie by (x-4) and will remained there 1/(x-4) than your exercise will be like resulte 1 1 ---------- = ---------- (x-4)(x-4) (x-4)^2

jhonyy9 (jhonyy9):

do you understand it sure ?

jhonyy9 (jhonyy9):

right ?

jhonyy9 (jhonyy9):

Tiffani is it ok ?

OpenStudy (anonymous):

yes thank you (:

jhonyy9 (jhonyy9):

yw good luck bye